Your division is considering two projects with the following cash flows (in millions):

0 1 2 3
Project A -$17 $8 $8 $3
Project B -$26 $13 $10 $9
  1. What are the projects' NPVs assuming the WACC is 5%? Enter your answer in millions. For example, an answer of $10,550,000 should be entered as 10.55. Negative values, if any, should be indicated by a minus sign. Do not round intermediate calculations. Round your answer to two decimal places.

    Project A:    $   million

    Project B:    $   million

    What are the projects' NPVs assuming the WACC is 10%? Enter your answer in millions. For example, an answer of $10,550,000 should be entered as 10.55. Negative values, if any, should be indicated by a minus sign. Do not round intermediate calculations. Round your answer to two decimal places.

    Project A:  $   million

    Project B:  $   million

    What are the projects' NPVs assuming the WACC is 15%? Enter your answer in millions. For example, an answer of $10,550,000 should be entered as 10.55. Negative values, if any, should be indicated by a minus sign. Do not round intermediate calculations. Round your answer to two decimal places.

    Project A:  $   million

    Project B:  $   million

  2. What are the projects' IRRs assuming the WACC is 5%? Do not round intermediate calculations. Round your answer to two decimal places.

    Project A:    %

    Project B:    %

    What are the projects' IRRs assuming the WACC is 10%? Do not round intermediate calculations. Round your answer to two decimal places.

    Project A:    %

    Project B:    %

    What are the projects' IRRs assuming the WACC is 15%? Do not round intermediate calculations. Round your answer to two decimal places.

    Project A:    %

    Project B:    %

  3. If the WACC was 5% and A and B were mutually exclusive, which project would you choose? (Hint: The crossover rate is 20.19%.)

    -Select-Project AProject BNeither A nor BItem 13

    If the WACC was 10% and A and B were mutually exclusive, which project would you choose? (Hint: The crossover rate is 20.19%.)

    -Select-Project AProject BNeither A nor BItem 14

    If the WACC was 15% and A and B were mutually exclusive, which project would you choose? (Hint: The crossover rate is 20.19%.)

    -Select-Project AProject BNeither A nor BItem 15
Answer #1

a. To calculate NPV, use NPV function in EXCEL

=NPV(rate,Year1 to year3 cashflows)-Year0cashflow

project A,NPV@5% WACC=NPV(5%,Year1 to year3 cashflows)-17=$0.47 million

Project B,NPV@5% WACC=NPV(5%,Year1 to year3 cashflows)-26=$3.23 million

project A,NPV@10% WACC=NPV(10%,Year1 to year3 cashflows)-17=-$0.86 million

Project B,NPV@10% WACC=NPV(10%,Year1 to year3 cashflows)-26=$0.84 million

project A,NPV@15% WACC=NPV(15%,Year1 to year3 cashflows)-17=-$2.02 million

Project B,NPV@15% WACC=NPV(15%,Year1 to year3 cashflows)-26=-$1.22 million

b. For IRR there is no need to use WACC but only take decision whether to accept teh project or not. If IRR>WACC, project should be accepted and if IRR<WACC, project needs to be rejected.

Use IRR function in EXCEL

=IRR(Year0 to Year3 cashflows)

Project A, IRR=IRR(Year0 to Year3 cashflows)=6.68%

Project B, IRR=IRR(Year0 to Year3 cashflows)=11.96%

Project A Project B
Year0 -17.0 -26.0
Year1 8.0 13.0
Year2 8.0 10.0
Year3 3.0 9.0
5% NPV 0.47 3.23
10% NPV -0.86 0.84
15% NPV -2.02 -1.22
IRR 6.68% 11.96%

c. If WACC is 5%, Project B has to be selected because  of higher NPV than Project A NPV

if WACC is 10%, Project B has to be selected because  of higher NPV than Project A NPV

If WACC is 15%, Both Projects A and B are need to be rejected becasue of negatiive NPV's